OpenStudy (anonymous):

If the current in a wire is at 0.5 A and the voltage is measured at 100 volts, what is the resistance in the wire? Show all work and include the proper units.

OpenStudy (kropot72):

\[R=\frac{V}{I}=\frac{100}{0.5}\ ohms\]

OpenStudy (anonymous):

R= V/I = 100/0.5= 200ohm